Red-green colorblindness is an X-linked, recessive trait. Suppose a group of 100 female and 100 male astronauts is sent to colonize a planet. 20 of the males are colorblind. 10 of the females are colorblind, and 30 are carriers. If this population were to establish a randomly mating colony in regards to this trait, what would you predict the frequency of colorblind males to be in the population over time if it were to immediately achieve and remain in Hardy-Weinberg equilibrium?
Answer #1

Total male = 100

Total female = 100

Color blind male (XcY) = 20

Normal male (XY) = 80

Color blind female (XcXc) = 10

Carrier females (XcX) = 30

Normal female (XX) = 60

Frequency of an allele in male population = number of alleles in males / total number of alleles in males

Frequency of X allele in male = 80 / 100x2 = 80/200 = 0.4

Frequency of Y allele in male = 100 / + 100x2 = 100/200 = 0.5

Frequency of Xc allele in male = 20 / 100x2 = 20/200 = 0.1

Frequency of X allele in female= 60x2 +30 / 100x2 = 150/200 = 0.75

Frequency of Xc allele in female = 10x2 + 30/100x2 = 50/200 = 0.25

Probability of X allele of male mating X allele of female = 0.4 x 0.75 = 0.3 XX

Probability of X allele of male mating Xc allele of female = 0.4 x 0.25 = 0.1 XXc

Probability of Y allele of male mating X allele of female = 0.5 x 0.75 = 0.375 XY

Probability of Y allele of male mating Xc allele of female = 0.5 x 0.25 = 0.125 XcY

Probability of Xc allele of male mating X allele of female = 0.1 x 0.75 = 0.075 XXc

Probability of Xc allele of male mating Xc allele of female = 0.1 x 0.25 = 0.025 XcXc

frequency of color blind males (Xc Y) is 0.125 ie 12.5% of males are color blind